# A-SDM: Accelerating Stable Diffusion through Model Assembly and Feature Inheritance Strategies

### v1

@@ Abstract
journal: IEEE Transactions on Neural Networks and Learning Systems
{CJK}
UTF8gbsn
1 Introduction
@@
For the second issue, our goal is to develop a tuning-free method to substitute the distillation method, enabling the skipping of trivial layer calculations, as illustrated in Fig 2 (b). Given that the UNet performs iterative denoising processes, the features in adjacent UNet inference processes exhibit similarity. Leveraging the observation, we propose a feature inheritance strategy to skip insignificant calculations in the current step by inheriting features from the previous step. Furthermore, considering the characteristics of UNet components, we explore local skip designs for shallow layers, deep layers, encoders, and decoders, respectively. Additionally, we investigate multiple sampling modes for feature inheritance at the time-step level, to determine the optimal approach for enhancing inference efficiency while preserving model performance.
@@ 2.6 Untrained Acceleration
3 Preliminary
3.1 Diffusion Principle
@@ 3.2 UNet Architecture
4 Model Assembly Strategy
4.1 Model Compression and Reconstruction
@@
4.2 Multi-Expert Conditional Convolution
@@
4.3 Multi-UNet Switching Approach
T2I task aims to generate target images from noise, whereas early-stage images of the denoising process tend to be relatively rough and lack detailed information. Conversely, the later stage requires polishing to enhance image quality. This insight motivates us to design multiple UNets for a 25-step T2I process. Initially, a compressed UNet (Base) is employed to rapidly generate the prototype in the initial stage (first 10 steps). Subsequently, the original UNet is utilized in the later stage (last 15 steps) for image optimization. The experimental results in Tab. 3 demonstrate that this strategy yields the best performance. It enhances speed and carries implications for our subsequent feature inheritance strategy of Sec. 5.5.
5 Feature Inheritance Strategy
Fig. 2 (b) depicts a straightforward network acceleration scheme where the computation of one layer per block in the UNet is uniformly reduced. In Sec. 4, we proposed a tuning strategy to distill the layer-pruned compressed model. In this section, we introduce an tuning-free feature inheritance strategy designed to skip certain layer calculations based on the circular inference mechanism of diffusion models.
@@ 5 Feature Inheritance Strategy
5.1 Layer-level Feature Inheritance
We first leverage the feature inheritance strategy to achieve layer-level computation skip, corresponding to the layers distilled away in Sec. 4. Specifically, we skip the second layer in down blocks and the middle layer in up blocks, as depicted in Fig. 7 (c). Fig. 7 (a) shows the simplest interval inheritance method with a period of 2 steps, performing a layer-level inheritance UNet (c) inference after a conventional UNet inference (b), and then repeating the above process. We extract the features (dotted green arrow) of the residual branch from the ResUnit and AttnUnit in the key layer from the previous step and store these features in a storage center. Subsequently, these stored features are inherited (solid green arrow) to the skip layer (white block in (c)) in the current step to accelerate.
@@ 5.1 Layer-level Feature Inheritance
5.2 Unit-level Feature Inheritance
Layer-level feature inheritance involves inheriting features for both AttnUnit and ResUnit of a certain layer, whereas unit-level feature inheritance entails inheriting features for either AttnUnit or ResUnit. Fig. 8 (a) and (b) illustrate the feature inheritance for AttnUnit and ResUnit, respectively. This section investigates the role of AttnUnit and ResUnit in the SDM denoising process. Similar to layer-level feature inheritance, the sampling mode with period of 2 is adopted in the unit-level feature inheritance, and features are inherited only for units in down and up blocks.
@@ 5.2 Unit-level Feature Inheritance
5.3 Block-level Feature Inheritance
@@
5.4 Concurrent Feature Inheritance
@@
5.5 Sampling Mode Design
@@
In this section, we delve into the sampling mode of feature inheritance. Feature inheritance includes extraction operation and inheritance operation. As depicted in Fig. 11 (c), the $t$ denoising step (green part) executes the extraction operation, while the inheritance operation (white part) occurs in steps $t+1$, $t+2$, $t+3$, and $t+4$ steps, followed by repetition. The entire sampling process consists of 50 steps, defining the above mode as $P5$ (feature inheritance period of 5). Informed by the findings of Sec. 4.3, where employing a small network in the early denoising stage and a complete UNet in the later stage often yields superb generation results, we introduce the specially designed $P5{\dagger}$ sampling mode (Fig. 11 (a)). Here, the last 10 steps perform conventional denoising operations, while periodic feature inheritance is employed in the early stage, akin to $P5$. Similarly, as illustrated in Fig. 11 (b), we devise the $P2{\dagger}$ sampling mode with a period of 2.
@@
6 Experiments
6.1 Datasets and Metrics
@@ 6.3 Experimental Analysis
ME-CondConv study: We seek to enhance the capacity of the compact models (Base, Small, Tiny) proposed by BK-SDM using ME-CondConv without increasing computational latency. Tab. 2 illustrates the replacement of $3\times 3$ convolution in all compact models with ME-CondConv. Notably, when employing 2 experts, Small with ME-CondConv and Tiny with ME-CondConv exhibit significant score increases, while Base with ME-CondConv shows a decline in performance. This disparity can be attributed to the extensive pruning of deep modules in Small and Tiny models, necessitating capacity augmentation through additional experts, unlike the Base model, which does not face such constraints.
@@
Table 2: Comparative experiment with or without ME-CondConv. The number of experts in ME-CondConv is set to 2. The weight of the first expert is inherited from the convolution weight of the teacher UNet (Runway 1.5 version), while the weight of the other expert is randomly initialized. Base, Small, and Tiny are the models provided by BK-SDM [34]. The sampling step is 25. Red marks indicate the best results.
@@
Table 3: Experiments of the multi-UNet switching method. In Strategy 1 (S1), the UNet of the Base model is adopted for the first 10 steps, followed by the UNet of the standard SDM for the next 15 steps. S2 and S3 is the opposite. Red marks indicate the best results in S1, S2, and S3.
@@
Multi-UNet switching study: Tab. 3 presents the experimental outcomes of multi-UNet switching processing. In the T2I task, the initial input is Gaussian noise, and the generated image is acquired through the cyclic process of UNet de-noising. Initially, the input is a chaotic, low-information, noise-like image, making a compact network (Base) efficient for processing. However, as the process advances, the input image begins to exhibit a preliminary semantic structure, necessitating a high-performance UNet (SDM) to understand and optimize it comprehensively. In Tab. 3, Strategy 1 (S1) employs a compressed model (Base) for the initial processing of 10 steps in the early stage, followed by SDM for optimization over 15 steps in the later stage. Conversely, Strategies 2 and 3 (S2, S3) have the opposite setup for comparative experiments. The experimental comparison reveals that the S1 strategy exhibits commendable performance, being both efficient and rapid. This conclusion is also extended to tuning-free feature inheritance strategies.
@@
Quantitative and qualitative analysis: Tab. 4 illustrates the performance of the proposed method (M2) compared to other methods. The proposed reconstructed model M2, obtained through the model assembly strategy, exhibits clear advantages in terms of generation score compared to most methods, including BK-SDM [34], SDM [1], and SnapFusion [35]. Moreover, the proposed model significantly reduces the number of parameters compared to SDM. Despite having more parameters than the BK-SDM-Base model, M2 still shows significantly improved speed due to pruning for shallow layers of computational redundancy. The single image generation speeds (25 steps) of SDM, BK-SDM-Base, and M2 on a 2080Ti GPU are 2.128s, 1.529s, and 1.643s, respectively. Considering both speed and performance comprehensively, the M2 model is superior.
It is worth noting that the required training resources are comparable to those of previous work [34]. For uniform comparison, the training time is measured by A100 days in Tab. 4. In practice, we utilized 8 V100s to train the M2 model on equivalent conditions. In the first step of model assembly, as in [34], the batch size is set to $512$ and the iteration to 25K. The setup for the second step of model assembly is identical, resulting in a total of 50K iterations. Since the second step involves partial freezing, the training time is less than 13 A100 days, with a total training time of under 26 A100 days.
In summary, the model assembly strategy enables the acquisition of an efficient generation model with improved performance, reduced parameters, and faster speed, all within the constraints of limited data (0.22M), low computational requirements (single A100 GPU), and short training time (less than 26 A100 days). Fig. 12 compares the generated results of the models (M1, M2) obtained by the model assembly strategy with the output results of the teacher model SDM. Refer to B for more visualizations, which comprehensively compare SDM, M1, M2, Base, Small, and Tiny models.
@@
Feature inheritance structure study: As displayed in Tab. 5, we compare different feature inheritance structures using the $P5{\dagger}$ sampling mode. Layer-level feature inheritance LI1, which skips one layer per block, demonstrates superior performance and efficient speedup. Structurally, when the feature inheritance operation area is reduced (LI2 and LI3 in Fig. 10), the performance on the FID metric deteriorates.
@@
Sampling mode study: After comparing the differences in feature inheritance structures, we further delved into the impact of sampling modes. Tab. 5 compares the outcomes of $P5{\dagger}$ and $P5$, revealing that the results of $P5{\dagger}$, which excludes feature inheritance in the last 10 steps, generally outperform those of $P5$, which employs feature inheritance throughout the process. This aligns with the conclusion drawn in Sec. 4.3, emphasizing the importance of fully optimizing the final stage of the inference process with an original UNet to enhance generation quality.
Comparing CO3, CO5, CO6, LI1, and RI in $P5{\dagger}$ and $P2{\dagger}$, we observe that extending the feature inheritance period to 5 leads to a rapid reduction in the FID score and an acceleration in network computing speed. This suggests that the shallow layer responsible for detail optimization has a significant impact on the FID index. The superior-performing strategies in Tab. 5 typically skip the deep and encoder parts of the calculation, focusing on iteratively optimizing the shallow and decoder parts, particularly the attention units in these parts. However, with a longer feature inheritance period, the calculation of the AttnUnit part across the entire network diminishes, leading to a slight decline in the CLIP index. Additionally, due to the reduction in the deep part calculation, the IS metric also exhibits a moderate decline. We show the results of sampling mode $Pn$ ($n$ represents period) for different inheritance periods under the CO6 structure in C , and it can be seen that the generation quality starts to decline as the period grows to 10.
@@ 7 Conclusion
Appendix A Additional Results of A-SDM
@@
Appendix B Additional Results of Model Assembly Strategy
@@
Appendix C Additional Results of Feature Inheritance Strategy
Figure 16: Additional results of concurrent feature inheritance strategy CO6 under different feature inheritance periods ($P2$-$P15$). In $Pn$ sampling mode, when the period $n$ of feature inheritance is extended to 8-15, the image quality is significantly reduced.

### v3

@@ Abstract
I Introduction
@@
For the second issue, our goal is to develop a tuning-free method to substitute the distillation method, enabling the skipping of trivial layer calculations, as shown in Fig 2 (b). Given that the UNet performs iterative denoising processes, the features in adjacent UNet inference processes exhibit similarity. Leveraging the observation, we propose a feature inheritance strategy to skip insignificant calculations in the current step by inheriting features from the previous step. Furthermore, considering the characteristics of UNet components, we explore local skip designs for shallow layers, deep layers, encoders, and decoders, respectively. Besides, we investigate multiple sampling modes for feature inheritance at the time-step level, to determine the optimal approach for enhancing inference efficiency while preserving model performance.
@@ II-F Untrained Acceleration
III Preliminary
III-A Diffusion Principle
@@ III-B UNet Architecture
IV Model Assembly Strategy
IV-A Model Compression and Reconstruction
@@
IV-B Multi-Expert Conditional Convolution
@@
IV-C Multi-UNet Switching Approach
T2I task aims to generate target images from noise, whereas early-stage images of the denoising process tend to be relatively rough and lack detailed information. Conversely, the later stage requires polishing to enhance image quality. This insight motivates us to design multiple UNets for a 25-step T2I process. Initially, a compressed UNet (Base) is employed to rapidly generate the prototype in the initial stage (first 10 steps). Subsequently, the original UNet is utilized in the later stage (last 15 steps) for image optimization. The experimental results in Tab. III demonstrate that this strategy yields the best performance. It enhances speed and carries implications for our subsequent feature inheritance strategy of Sec. V-E.
V Feature Inheritance Strategy
Fig. 2 (b) depicts a straightforward network acceleration scheme where the computation of one layer per block in the UNet is uniformly reduced. In Sec. IV, we proposed a tuning strategy to distill the layer-pruned compressed model. In this section, we introduce an tuning-free feature inheritance strategy designed to skip certain layer calculations based on the circular inference mechanism of diffusion models.
@@ V Feature Inheritance Strategy
V-A Layer-level Feature Inheritance
We first leverage the feature inheritance strategy to achieve layer-level computation skip, corresponding to the layers distilled away in Sec. IV. Specifically, we skip the second layer in down blocks and the middle layer in up blocks, as depicted in Fig. 7 (c). Fig. 7 (a) shows the simplest interval inheritance method with a period of 2 steps, performing a layer-level inheritance UNet (c) inference after a conventional UNet inference (b), and then repeating the above process. We extract the features (dotted green arrow) of the residual branch from the ResUnit and AttnUnit in the key layer from the previous step and store these features in a storage center. Subsequently, these stored features are inherited (solid green arrow) to the skip layer (white block in (c)) in the current step to accelerate.
@@ V-A Layer-level Feature Inheritance
V-B Unit-level Feature Inheritance
Layer-level feature inheritance involves inheriting features for both AttnUnit and ResUnit of a certain layer, whereas unit-level feature inheritance entails inheriting features for either AttnUnit or ResUnit. Fig. 8 (a) and (b) illustrate the feature inheritance for AttnUnit and ResUnit, respectively. This section investigates the role of AttnUnit and ResUnit in the SDM denoising process. Similar to layer-level feature inheritance, the sampling mode with a period of 2 is adopted in the unit-level feature inheritance, and features are inherited only for units in down and up blocks.
@@ V-B Unit-level Feature Inheritance
V-C Block-level Feature Inheritance
@@
V-D Concurrent Feature Inheritance
@@
V-E Sampling Mode Design
@@
In this section, we delve into the sampling mode of feature inheritance. Feature inheritance includes extraction operation and inheritance operation. As depicted in Fig. 11 (c), the $t$ denoising step (green part) executes the extraction operation, while the inheritance operation (white part) occurs in steps $t+1$, $t+2$, $t+3$, and $t+4$ steps, followed by repetition. The entire sampling process consists of 50 steps, defining the above mode as $P5$ (feature inheritance period of 5). Informed by the findings of Sec. IV-C, where employing a small network in the early denoising stage and a complete UNet in the later stage often yields superb generation results, we introduce the specially designed $P5{\dagger}$ sampling mode (Fig. 11 (a)). Here, the last 10 steps perform conventional denoising operations, while periodic feature inheritance is employed in the early stage, akin to $P5$. Similarly, as illustrated in Fig. 11 (b), we devise the $P2{\dagger}$ sampling mode with a period of 2.
VI Experiments
VI-A Datasets and Metrics
@@ VI-C Experimental Analysis
ME-CondConv study: We seek to enhance the capacity of the compact models (Base, Small, Tiny) proposed by BK-SDM using ME-CondConv without increasing computational latency. Tab. II illustrates the replacement of $3\times 3$ convolution in all compact models with ME-CondConv. Notably, when employing 2 experts, Small with ME-CondConv and Tiny with ME-CondConv exhibit significant score increases, while Base with ME-CondConv shows a decline in performance. This disparity can be attributed to the extensive pruning of deep modules in Small and Tiny models, necessitating capacity augmentation through additional experts, unlike the Base model, which does not face such constraints.
@@
TABLE II: Comparative experiment with or without ME-CondConv. The number of experts in ME-CondConv is set to 2. The weight of the first expert is inherited from the convolution weight of the teacher UNet (Runway 1.5 version), while the weight of the other expert is randomly initialized. Base, Small, and Tiny are the models provided by BK-SDM [34]. The sampling step is 25. Red marks indicate the best results.
@@
TABLE III: Experiments of the multi-UNet switching method. In Strategy 1 (S1), the UNet of the Base model is adopted for the first 10 steps, followed by the UNet of the standard SDM for the next 15 steps. S2 and S3 is the opposite. Red marks indicate the best results in S1, S2, and S3.
@@
Multi-UNet switching study: Tab. III presents the experimental outcomes of multi-UNet switching processing. In the T2I task, the initial input is Gaussian noise, and the generated image is acquired through the cyclic process of UNet de-noising. Initially, the input is a chaotic, low-information, noise-like image, making a compact network (Base) efficient for processing. However, as the process advances, the input image begins to exhibit a preliminary semantic structure, necessitating a high-performance UNet (SDM) to understand and optimize it comprehensively. In Tab. III, Strategy 1 (S1) employs a compressed model (Base) for the initial processing of 10 steps in the early stage, followed by SDM for optimization over 15 steps in the later stage. Conversely, Strategies 2 and 3 (S2, S3) have the opposite setup for comparative experiments. The experimental comparison reveals that the S1 strategy exhibits commendable performance, being both efficient and rapid. This conclusion is also extended to tuning-free feature inheritance strategies.
@@
Quantitative and qualitative analysis: Tab. IV illustrates the performance of the proposed method (M2) compared to other methods. The proposed reconstructed model M2, obtained through the model assembly strategy, exhibits clear advantages in terms of generation score compared to most methods, including BK-SDM [34], SDM [1], and SnapFusion [35]. Moreover, the proposed model significantly reduces the number of parameters compared to SDM. Despite having more parameters than the BK-SDM-Base model, M2 still shows significantly improved speed due to pruning for shallow layers of computational redundancy. The single image generation speeds (25 steps) of SDM, BK-SDM-Base, and M2 on a 2080Ti GPU are 2.128s, 1.529s, and 1.643s, respectively. Considering both speed and performance comprehensively, the M2 model is superior.
It is worth noting that the required training resources are comparable to those of previous work [34]. For uniform comparison, the training time is measured by A100 days in Tab. IV. In practice, we utilized 8 V100s to train the M2 model on equivalent conditions. In the first step of model assembly, as in [34], the batch size is set to $512$ and the iteration to 25K. The setup for the second step of model assembly is identical, resulting in a total of 50K iterations. Since the second step involves partial freezing, the training time is less than 13 A100 days, with a total training time of under 26 A100 days.
In summary, the model assembly strategy enables the acquisition of an efficient generation model with improved performance, reduced parameters, and faster speed, all within the constraints of limited data (0.22M), low computational requirements (single A100 GPU), and short training time (less than 26 A100 days). Fig. 12 compares the generated results of the models (M1, M2) obtained by the model assembly strategy with the output results of the teacher model SDM. Refer to Appendix B. Additional Results of Model Assembly Strategy for more visualizations, which comprehensively compare SDM, M1, M2, Base, Small, and Tiny models.
@@
Feature inheritance structure study: As displayed in Tab. V, we compare different feature inheritance structures using the $P5{\dagger}$ sampling mode. Layer-level feature inheritance LI1, which skips one layer per block, demonstrates superior performance and efficient speedup. Structurally, when the feature inheritance operation area is reduced (LI2 and LI3 in Fig. 10), the performance on the FID metric deteriorates.
@@
Sampling mode study: After comparing the differences in feature inheritance structures, we further delved into the impact of sampling modes. Tab. V compares the outcomes of $P5{\dagger}$ and $P5$, revealing that the results of $P5{\dagger}$, which excludes feature inheritance in the last 10 steps, generally outperform those of $P5$, which employs feature inheritance throughout the process. This aligns with the conclusion drawn in Sec. IV-C, emphasizing the importance of fully optimizing the final stage of the inference process with an original UNet to enhance generation quality.
Comparing CO3, CO5, CO6, LI1, and RI in $P5{\dagger}$ and $P2{\dagger}$, we observe that extending the feature inheritance period to 5 leads to a rapid reduction in the FID score and an acceleration in network computing speed. This suggests that the shallow layer responsible for detail optimization has a significant impact on the FID index. The superior-performing strategies in Tab. V typically skip the deep and encoder parts of the calculation, focusing on iteratively optimizing the shallow and decoder parts, particularly the attention units in these parts. However, with a longer feature inheritance period, the calculation of the AttnUnit part across the entire network diminishes, leading to a slight decline in the CLIP index. Additionally, due to the reduction in the deep part calculation, the IS metric also exhibits a moderate decline. We show the results of sampling mode $Pn$ ($n$ represents period) for different inheritance periods under the CO6 structure in Appendix C. Additional Results of Feature Inheritance Strategy , and it can be seen that the generation quality starts to decline as the period grows to 10.
@@ VII Conclusion
Appendix A. Additional Results of A-SDM
@@
Appendix B. Additional Results of Model Assembly Strategy
@@
Appendix C. Additional Results of Feature Inheritance Strategy
Figure 16: Additional results of concurrent feature inheritance strategy CO6 under different feature inheritance periods ($P2$-$P15$). In the $Pn$ sampling mode, when the period $n$ of feature inheritance is extended to 8-15, the image quality is significantly reduced.